# ORCA: An Open-Source, Reliable, Cost-Effective, Anthropomorphic Robotic Hand for Uninterrupted Dexterous Task Learning

### v1

@@ Abstract
General-purpose robots should possess human-like dexterity and agility to perform tasks with the same versatility as us. A human-like form factor further enables the use of vast datasets of human-hand interactions. However, the primary bottleneck in dexterous manipulation lies not only in software but arguably even more in hardware. Robotic hands that approach human capabilities are often prohibitively expensive, bulky, or require enterprise-level maintenance, limiting their accessibility for broader research and practical applications. What if the research community could get started with reliable dexterous hands within a day? We present the open-source ORCA hand, a reliable and anthropomorphic 17-DoF tendon-driven robotic hand with integrated tactile sensors, fully assembled in less than eight hours and built for a material cost below 2,000 CHF. We showcase ORCA’s key design features such as popping joints, auto-calibration, and tensioning systems that significantly reduce complexity while increasing reliability, accuracy, and robustness. We benchmark the ORCA hand across a variety of tasks, ranging from teleoperation and imitation learning to zero-shot sim-to-real reinforcement learning. Furthermore, we demonstrate its durability, withstanding more than 10,000 continuous operation cycles–equivalent to approximately 20 hours–without hardware failure, the only constraint being the duration of the experiment itself. All design files, source code, and documentation will be available at orcahand.com.
@@ II System Design
The ORCA hand design follows the requirements set above, namely dexterity and reliability at minimal complexity and cost. It comprises five fingers, including an opposable thumb and an actuated wrist (Figure 1-A), and is of a size similar to the average human hand [21]. The fingers are mounted on a base that resembles the human palm containing the carpal and metacarpal bones. The palm is connected to the wrist mechanism that is mounted on the tower. The tower contains the motors and all auxiliary electronics, and is enclosed in a protective casing. The rest of this section describes the most important design features of the ORCA hand in detail.
@@ II-A Tendon Actuation for Agility
  • We avoid direct contact of the tendons with polylactic acid (PLA) by deflecting the tendons around smooth metal pins and rods, as depicted in Figure 3-C.
@@
  • Finally, tendons can be manually retensioned using a ratchet spool mechanism mounted on the motors, as shown in Figure 1-A3. The ratchet is attached to the top spool, allowing rotation in one direction while locking movement in the other. This design makes the ORCA hand user-friendly, as re-tensioning can be done in seconds without the need to unscrew the spool or tendon. The spool system quickly removes any slack that accumulates. In addition, tendons can be easily loosened, allowing joints to pop out for quick replacement of broken parts.
@@ II-B Poppable Pin Joints
Rolling contact joints [22, 18] have become a popular alternative to pinhole joints for 3D-printed hands due to their ability to dislocate instead of break. However, these mechanisms require ligaments that can potentially loosen over time and increase complexity. We introduce a pin joint design that allows the joints to "pop" out of place and dislocate instead of breaking when excessive radial and axial loads are applied (Figure 1-A1).
@@ II-C Finger and Palm Design
Fingers two to five (index to pinky) have three actuated joints that mimic those of the human finger: the Proximal Interphalangeal (PIP), the Metacarpophalangeal (MCP), and the Abduction (ABD) joint (Figure 3-A). The ORCA PIP joint corresponds to the human PIP joint, while the MCP and ABD joints together correspond to the human MCP joint which can perform both flexion/extension and abduction/adduction, which is necessary in various dexterous manipulation tasks [23]. The range of motion (RoM) for each joint is shown in Table I, based on human anatomy [21].
@@
The removal of the DIP joints on the digits 2 to 5 was an intentional choice with the benefit of greatly simplifying assembly and reliability. The DIP joint is not directly actuated in the human hand, and as such it is of less importance compared to the other hand joints in dexterous manipulation tasks. A positive side-effect of the removal of the DIP joint is the substantial increase in the space available for tactile sensing integration on the fingertips.
@@ II-F Self-Calibration for Accurate Control
In machine learning settings, such as data collection for imitation learning, consistency in recordings is critical. If there is a high variance in the gap between human teleoperation and the robotic hand in different runs, it becomes more difficult for policies to accurately mimic the human [29]. Moreover, this consistency is crucial not only during data collection, but also during policy rollouts. Any offset between training data and the deployed system can result in poor policy performance, as the learned model may not be able to generalize effectively to scenarios with mismatched conditions. This is especially important when attempting to reduce the sim-to-real gap. Although it is possible to ensure reliable control consistency using proprioception sensors, as demonstrated in [14], this adds significant cost and complexity. We therefore propose an alternative, more cost-effective approach:
Let $\theta_{i}$ be the true angle, $\tilde{\theta_{i}}$ the commanded angle of the joint $i$, $\forall i\in\{1,\dots,17\}$, and $\phi_{i}$ the angular displacement of the motor for the same joint. The ideal goal is to always have $\theta_{i}=\tilde{\theta_{i}}$. However, in practice, especially with tendon-driven hands, there is a discrepancy because motor-to-joint mapping relies on a model where:
${\theta_{i}}$ $=\tilde{\theta_{i}}+\epsilon(r_{j},r_{m},s,d,m)$
$=$ $f(\phi_{1},\dots,\phi_{17})+\epsilon(r_{j},r_{m},s,d,m)$
where $f(\phi_{1},\dots,\phi_{17})$ is generally nonlinear and $\epsilon$ represents model errors including tendon radius of joint and motor ($r_{j}$, $r_{m}$), build-up slack $s$, servo drift $d$ and measurement errors $m$ in the case of manual calibration which can introduce substantial offsets.
@@
Referring to the routing of the ORCA hand, as depicted in Figure 3, each tendon passes through or near the center of rotation (CoR). This design ensures that joint positions are approximately decoupled and can be actuated linearly and independently. Due to the linearity of the system and our focus on controlling the motor angles $\phi_{i}$, we can express the desired angles as $\tilde{\theta}_{i}=R_{i}^{\prime}\cdot\phi_{i}$, with $R_{i}^{\prime}\in\mathbb{R^{+}}$. Consequently, the relationship between the actual angles $\theta_{i}$ and the motor angles yields
$\theta_{i}=R_{i}^{\prime}\cdot\phi_{i}+\epsilon(r_{M},r_{J},s,d,m).$
Since we can directly control the motor angles, we define $\phi_{i}$ in terms of $\tilde{\theta}_{i}$ as $\phi_{i}=\frac{1}{R_{i}^{\prime}}\cdot\tilde{\theta}_{i}=R_{i}\cdot\tilde{\theta}_{i}$. To determine $R_{i}$, we use a procedure referred to as autocalibration, which estimates $R_{i}$ and minimizes the error terms. The process is visually depicted in Figure 3-B and works as follows:
(1) Let $|RoM_{i}|$ denote the absolute range of motion of the joint $i$. Using CAD models, the angular distance between the physical stops of each joint is precisely determined, providing $|RoM_{i}|$ with minimal uncertainty. (2) During calibration, each joint $i$ is moved to its extreme mechanical limits (fully flexed $\phi_{i,\text{max}}$, fully extended $\phi_{i,\text{min}}$). (3) The total servo rotation over the joint’s range of motion is computed as:
$\Delta\phi_{i}=\phi_{i,\text{max}}-\phi_{i,\text{min}}.$
@@
(4) The joint-to-motor transmission ratio is then calculated using linear interpolation:
$R_{i}=\frac{\Delta\phi_{i}}{|RoM_{i}|}.$
and thus we get as the true angle
${\theta_{i}}=\frac{1}{R_{i}}\cdot\phi_{i}+\epsilon(|RoM_{i}|,s,d)=R_{i}^{\prime}\cdot\phi_{i}+\epsilon(|RoM_{i}|,s,d)$
where $\epsilon(|RoM_{i}|,s,d)\ll\epsilon(r_{M},r_{J},s,d,m)$.
@@ III-A Reliability and Robustness
To evaluate the reliability and robustness of the ORCA hand in long-duration tasks, we conducted an experiment in which we actuate the hand joints continuously for 2.5 hours (Figure 1-C). We attach a plush animal to the palm of the hand and have it grasp it with all fingers every four seconds. This setup, using a compliant object, is similar to the repeatability test performed by [20] and allows us to assess behavior under increased stress over a greater range of motion. Moreover, to test the durability of the wrist joint, we flex and extend the wrist to $40^{\circ}$ at one fourth of the frequency of the finger, that is, every 16 seconds.
The hand reliably performs the same grasping movement for all 2,250 grasping cycles without breaking, motor shutdown, or excessive tendon slack buildup. In Figure 4-C, we display the maximum current in the motors actuating the middle finger’s MCP and PIP joints for each grasping cycle, as well as the maximum current of the wrist joint motor for each wrist extension-flexion cycle. The current used by each motor provides an estimate of how much external torque and friction must be overcome for a given angle change of the joint. The fact that the maximum current for each cycle remains roughly the same during 2.5 hours of uninterrupted joint movement demonstrates the robustness of the hand, the high repeatability of joint movement and its capacity for long-duration operation. The fans on the side of the hand’s tower prevent the motors from overheating (Figure 4-C), allowing the hand to function almost indefinitely. In particular, the experiment was not terminated due to failure, but was concluded after 2.5 hours because it was deemed a sufficient demonstration of the reliability of the system. More extended tests until hand failure, with the inclusion of sensor robustness evaluations, are planned for future work.
@@ III-B Accuracy and Latency
Figure 4: (A) Comparison of joint responses to sine wave signals at 2 Hz and 5 Hz for the ORCA and LEAP hands: Despite the tendon-driven design, the ORCA hand achieves similar accuracy and latency as the LEAP hand. (B) Experimental setup to benchmark accuracy and latency of ORCA and LEAP hand actuation responses: AprilTags are recorded with a camera and used to infer the ground truth joint angles, which are then synchronized with the commanded angles. (C) Reliability test: Uninterrupted and consistent object grasping (2200+ cycles) and wrist motion (550+ cycles) for 2.5 hours without any breakdown, motor overheating, or performance degradation, as seen in the motors’ unvarying maximum current and temperature curves.
@@
We compare the accuracy and latency of the ORCA hand with the LEAP [20] hand to benchmark our system’s tendon-driven dynamics against the dynamics of a direct-driven robot hand (Figure 4-A). We actuate the MCP and PIP joints of both hands’ index fingers with 2 Hz and 5 Hz sine wave patterns. Before each test, we leverage the auto-calibration mechanism of the ORCA hand to account for any changes in tendon length or slack that might have accumulated beforehand. We demonstrate that, through auto-calibration, the finger joints of the ORCA hand accurately follow the commanded sine input. In fact, we achieve similar accuracy to the LEAP hand while being far less bulky thanks to our tendon-driven actuation design. Additionally, we observe that the ORCA hand’s joints rotate more smoothly than those of the LEAP hand, which exhibit more jerks, presumably due to the cables and possibly increased inertia interfering with the LEAP hand’s finger joint motion. Both hands exhibit average latencies of less than 0.2 seconds, most of which comes from the software process. However, slack in the ORCA hand may introduce additional latency, which is why retensioning the spools periodically is important for robust performance.
One limitation of our benchmarking experiment is that we cannot evaluate the accuracy and latency of the hands for faster movements, as the AprilTags can not be tracked at sine wave frequencies below 2 Hz or for step signals due to motion blur at 60 fps. Recording the finger joints with a ROS2-compatible camera capable of 240 fps could enable more thorough system identification in the future.
@@ III-C Reinforcement Learning
Reinforcement learning (RL) is commonly used to learn dexterous tasks that are challenging to demonstrate or require fine motor control, such as object reorientation in the hand [30]. A key challenge in applying RL to dexterous manipulation tasks is that the policies learned in simulation often perform poorly when transferred to the real robotic hand. We use the IsaacGymEnvs wrapper from [22] to train 4096 ORCA hand models in parallel with an advantage actor-critic (A2C) architecture to learn in-hand ball reorientation. We demonstrate that after one hour of training with domain randomization, we can deploy a robust policy on the physical ORCA hand (Figure 6-B), that can successfully reorient a tennis ball along a given rotation axis.
@@ III-D Imitation Learning
Demonstrations were collected using motion capture gloves [10], which provided absolute wrist tracking and finger pose estimation. We retargeted these demonstrations into the robot’s state space using an energy-based minimization objective, similar to [36]. The wrist pose was used to control the robotic arm in cartesian end-effector space. This teleoperation method allowed us to showcase the versatility of the hand in a wide range of tasks (Figure 2) and facilitated the rapid and intuitive collection of demonstrations, even from non-trained operators.
@@
For the proposed task, we performed ablation studies on image pre-processing. Specifically, we compared three policy variations: (1) a baseline policy trained on raw RGB inputs, (2) a policy incorporating segmentation of the cube’s color, and (3) a hybrid approach trained on both data sets (Figure 7).
For (2) and (3), a binary mask was generated using the CIELAB  [37] color space, with parameters tuned to isolate the red cube. From the RGB image, a three-channel grayscale image is created and non-zero pixels of the binary mask are set to 255 in the first channel (arbitrarily chosen), effectively highlighting the cube.

### v2

@@ Abstract
General-purpose robots should possess human-like dexterity and agility to perform tasks with the same versatility as us. A human-like form factor further enables the use of vast datasets of human-hand interactions. However, the primary bottleneck in dexterous manipulation lies not only in software but arguably even more in hardware. Robotic hands that approach human capabilities are often prohibitively expensive, bulky, or require enterprise-level maintenance, limiting their accessibility for broader research and practical applications. What if the research community could get started with reliable dexterous hands within a day? We present the open-source ORCA hand, a reliable and anthropomorphic 17-DoF tendon-driven robotic hand with integrated tactile sensors, fully assembled in less than eight hours and built for a material cost below 2,000 CHF. We showcase ORCA’s key design features such as popping joints, auto-calibration, and tensioning systems that significantly reduce complexity while increasing reliability, accuracy, and robustness. We benchmark the ORCA hand across a variety of tasks, ranging from teleoperation and imitation learning to zero-shot sim-to-real reinforcement learning. Furthermore, we demonstrate its durability, withstanding more than 10,000 continuous operation cycles—equivalent to approximately 20 hours—without hardware failure, the only constraint being the duration of the experiment itself. Video is here: youtu.be/kUbPSYMmOds. Design files, source code, and documentation are available at srl.ethz.ch/orcahand.
@@ II System Design
The ORCA hand design follows the requirements set above, namely dexterity and reliability at minimal complexity and cost. It comprises five fingers, including an opposable thumb and an actuated wrist (Figure˜1-A), and is of a size similar to the average human hand [21]. The total weight of the hand is approximately 1.2 kg. The fingers are mounted on a base that resembles the human palm containing the carpal and metacarpal bones. The palm is connected to the wrist mechanism that is mounted on the tower. The tower contains the motors and all auxiliary electronics, and is enclosed in a protective casing. The rest of this section describes the most important design features of the ORCA hand in detail.
@@ II-A Tendon Actuation for Agility
  • We avoid direct contact of the tendons with polylactic acid (PLA) by deflecting the tendons around smooth metal pins and rods, as depicted in Figure˜3-C.
@@
  • Finally, tendons can be manually re-tensioned using a ratchet spool mechanism mounted on the motors, as shown in Figure˜1-A3. The ratchet is attached to the top spool, allowing rotation in one direction while locking movement in the other. This design makes the ORCA hand user-friendly, as re-tensioning can be done in seconds without the need to unscrew the spool or tendon. The spool system quickly removes any slack that accumulates. In addition, tendons can be easily loosened, allowing joints to pop out for quick replacement of broken parts.
@@ II-B Poppable Pin Joints
Rolling contact joints [22, 18] have become a popular alternative to pinhole joints for 3D-printed hands due to their ability to dislocate instead of break. However, these mechanisms require ligaments that can potentially loosen over time and increase complexity. We introduce a pin joint design that allows the joints to "pop" out of place and dislocate instead of breaking when excessive radial and axial loads are applied (Figure˜1-A1).
@@ II-C Finger and Palm Design
Fingers two to five (index to pinky) have three actuated joints that mimic those of the human finger: the Proximal Interphalangeal (PIP), the Metacarpophalangeal (MCP), and the Abduction (ABD) joint (Figure˜3-A). The ORCA PIP joint corresponds to the human PIP joint, while the MCP and ABD joints together correspond to the human MCP joint which can perform both flexion/extension and abduction/adduction, which is necessary in various dexterous manipulation tasks [23]. The range of motion (RoM) for each joint is shown in Table˜I, based on human anatomy [21].
@@
The removal of the DIP joints on digits 2 to 5 was an intentional design decision. In most cases, DIP joints are rigidly coupled with the corresponding PIP joints and do not move independently, meaning they do not contribute additional active degrees of freedom. By fixing the DIP joints, we eliminate the potential for slack buildup over time and enable direct motor control of the distal joint, resulting in improved accuracy and predictability. Furthermore, this approach simplifies assembly significantly and frees up space for additional sensing electronics. The DIP joint is not directly actuated in the human hand, and as such it is of less importance compared to the other hand joints in dexterous manipulation tasks. A positive side-effect of the removal of the DIP joint is the substantial increase in the space available for tactile sensing integration on the fingertips.
@@ II-F Self-Calibration for Accurate Control
Consistency across runs is essential in imitation learning, as variability between teleoperation and execution degrades policy performance [29], especially in sim-to-real settings. While proprioceptive sensors offer reliable control [14], they add cost and complexity. We propose a simpler, cost-effective alternative: automated self-calibration to estimate motor-to-joint mappings (Algorithm 1).
Algorithm 1 Self-Calibration Algorithm
1:Input: Set of joints $\mathcal{J}=\{1,\dots,N\}$, joint ROMs $\theta_{j}^{\min},\theta_{j}^{\max}]$ for each joint $j$
2:Output: Motor limits $m_{j}^{\min},m_{j}^{\max}]$, motor-to-joint ratios $\rho_{j}$ for each joint $j$
3:for each joint $j\in\mathcal{J}$ do
4:  Move joint $j$ towards its flexion limit until stall is detected
5:  $m_{j}^{\min}\leftarrow\text{current motor position}$
6:
7:  Move joint $j$ towards its extension limit until stall is detected
8:  $m_{j}^{\max}\leftarrow\text{current motor position}$
9:
10:  Compute motor travel: $\Delta m_{j}=m_{j}^{\max}-m_{j}^{\min}$
11:  Compute joint ROM: $\Delta\theta_{j}=\theta_{j}^{\max}-\theta_{j}^{\min}$
12:  Compute motor-to-joint ratio: $\rho_{j}=\frac{\Delta m_{j}}{\Delta\theta_{j}}$
13:end for
Let $\theta_{j}$ be the true angle and $m_{j}$ the position of the motor for joint $j$. The ideal goal is to perfectly control $\theta_{j}$ by commanding $m_{j}$. However, in practice, especially with tendon-driven hands without joint sensors, there is a discrepancy. A simple model would be $\theta_{j}=f(m_{1},\dots,m_{17})+\epsilon$, where $f$ is a nonlinear function and $\epsilon$ represents model errors arising from factors like tendon and pulley radii ($r_{t},r_{p}$), tendon slack $s$, servo drift $d$, and manual measurement errors $m_{err}$. These factors can introduce substantial offsets.
@@
Referring to the routing of the ORCA hand, as depicted in Figure˜3, each tendon passes through or near the center of rotation (CoR). This design ensures that joint positions are approximately decoupled and can be actuated linearly and independently. This linearity allows us to simplify the model significantly. The self-calibration procedure detailed in Algorithm˜1 leverages this by empirically finding the linear relationship for each joint. The process works as follows:
  1. 1.
    Determine Joint ROM (Line 11): The absolute range of motion (ROM) for each joint, $\Delta\theta_{j}$, is precisely determined from the robot’s CAD models. This provides an accurate ground truth for the joint’s angular travel between its physical stops, $\theta_{j}^{\min},\theta_{j}^{\max}]$.
  2. 2.
    Find Motor Limits (Lines 4-8): During calibration, each joint $j$ is programmatically moved to its extreme mechanical limits (e.g., fully flexed and fully extended). The corresponding motor positions, $m_{j}^{\min}$ and $m_{j}^{\max}$, are recorded when the joint stalls.
  3. 3.
@@
    Calculate Transmission Ratio (Lines 10-12): The total motor travel, $\Delta m_{j}$, is computed from the recorded limits. The motor-to-joint transmission ratio $\rho_{j}$ is then calculated via linear interpolation as the ratio of the motor travel to the joint’s known ROM:
    $\rho_{j}=\frac{\Delta m_{j}}{\Delta\theta_{j}}=\frac{m_{j}^{\max}-m_{j}^{\min}}{\theta_{j}^{\max}-\theta_{j}^{\min}}$
This autocalibration largely eliminates the complex error terms related to physical modeling and manual measurements, $\epsilon(r_{t},r_{p},s,d,m_{err})$, leaving a much smaller residual error, $\epsilon(\Delta\theta_{j},s,d)$. With the calibrated parameters, the motor position $m_{j}$ required to achieve a desired joint angle $\theta_{j}$ is computed using a linear mapping:
$m_{j}=m_{j}^{\min}+\rho_{j}\cdot(\theta_{j}-\theta_{j}^{\min})$
This method ensures a consistent and accurate mapping from desired joint angles to motor commands across different operational sessions.
@@ III-A Reliability and Robustness
To evaluate the reliability and robustness of the ORCA hand in long-duration tasks, we conducted an experiment in which we actuate the hand joints continuously for 2.5 hours (Figure˜1-C). We attach a plush animal to the palm of the hand and have it grasp it with all fingers every four seconds. This setup, using a compliant object, is similar to the repeatability test performed by [20] and allows us to assess behavior under increased stress over a greater range of motion. Moreover, to test the durability of the wrist joint, we flex and extend the wrist to $40^{\circ}$ at one fourth of the frequency of the finger, that is, every 16 seconds.
The hand successfully completed 2,250 grasping cycles without breakage, motor shutdown, or excessive tendon slack. Figure˜4-C shows the maximum current per cycle for the middle finger’s MCP and PIP motors, and the wrist joint motor during wrist extension-flexion. Motor current reflects the torque and friction the system must overcome. The stable maximum current over 2.5 hours of continuous operation demonstrates the hand’s robustness, high repeatability, and long-duration capability. Side-mounted fans (Figure˜4-C) prevent motor overheating, enabling near-continuous use. The experiment was ended voluntarily after 2.5 hours, not due to failure.
To evaluate payload capacity, we used a hand force sensor (electronic grip strength tester) with constant $600\text{\,}\mathrm{m}\mathrm{A}$ motor current in position control mode. The hand was tested in various grasps, showing it can hold up to $10.5\text{\,}\mathrm{k}\mathrm{g}$ ($103\text{\,}\mathrm{N}$) using all four fingers, and up to $2\text{\,}\mathrm{k}\mathrm{g}$ ($19.6\text{\,}\mathrm{N}$) using only the index finger.
@@ III-B Accuracy and Latency
Figure 4: (A) Joint response comparison at 0.2 Hz and 0.5 Hz: ORCA achieves accuracy and latency comparable to the LEAP hand despite its tendon-driven design. (B) Accuracy benchmarking setup using AprilTags to infer ground-truth joint angles synchronized with commands. (C) Reliability test: 2.5 hours of continuous grasping (2200+ cycles) and wrist motion (550+ cycles) without failure, overheating, or performance drop, as shown by stable motor current and temperature.
@@
We compare the accuracy and latency of the ORCA hand with the LEAP [20] hand to benchmark our system’s tendon-driven dynamics against the dynamics of a direct-driven robot hand (Figure˜4-A). We actuate the MCP and PIP joints of both hands’ index fingers with 0.2 Hz and 0.5 Hz sine wave patterns. Before each test, we leverage the auto-calibration mechanism of the ORCA hand to account for any changes in tendon length or slack that might have accumulated beforehand. We demonstrate that, through auto-calibration, the finger joints of the ORCA hand accurately follow the commanded sine input. In fact, we achieve similar accuracy to the LEAP hand while being far less bulky thanks to our tendon-driven actuation design. Additionally, we observe that the ORCA hand’s joints rotate more smoothly than those of the LEAP hand, which exhibit more jerks, presumably due to the cables and possibly increased inertia interfering with the LEAP hand’s finger joint motion. Both hands exhibit average latencies of less than 0.2 seconds, most of which comes from the software process. However, slack in the ORCA hand may introduce additional latency, which is why re-tensioning the spools periodically is important for robust performance.
One limitation of our benchmarking experiment is that we cannot evaluate the accuracy and latency of the hands for faster movements, as the AprilTags can not be tracked at sine wave frequencies above 0.5 Hz or for step signals due to motion blur at 60 fps. Recording the finger joints with a ROS2-compatible camera capable of 240 fps could enable more thorough system identification in the future.
@@ III-C Reinforcement Learning
Reinforcement learning (RL) is commonly used to learn dexterous tasks that are challenging to demonstrate or require fine motor control, such as object reorientation in the hand [30]. A key challenge in applying RL to dexterous manipulation tasks is that the policies learned in simulation often perform poorly when transferred to the real robotic hand. We use the IsaacGymEnvs wrapper from [22] to train 4096 ORCA hand models in parallel with an advantage actor-critic (A2C) architecture to learn in-hand ball reorientation. We do not use tactile sensors in our RL experiments, primarily due to the additional complexity involved in accurately modeling them. We demonstrate that after one hour of training with domain randomization, we can deploy a robust policy on the physical ORCA hand (Figure˜6-B), that can successfully reorient a tennis ball along a given rotation axis.
@@ III-D Imitation Learning
Demonstrations were collected using motion capture gloves [10], which provided absolute wrist tracking and finger pose estimation. We retargeted these demonstrations into the robot’s state space using an energy-based minimization objective, similar to [36]. The wrist pose was used to control the robotic arm in cartesian end-effector space. This teleoperation method allowed us to showcase the versatility of the hand in a wide range of tasks (Figure˜2) and facilitated the rapid and intuitive collection of demonstrations, even from non-trained operators.
@@
For the proposed task, we performed ablation studies on image pre-processing. Specifically, we compared three policy variations: (1) a baseline policy trained on raw RGB inputs, (2) a policy incorporating segmentation of the cube’s color, and (3) a hybrid approach trained on both data sets (Figure˜7).
For (2) and (3), a binary mask was generated using the CIELAB  [37] color space, with parameters tuned to isolate the red cube. From the RGB image, a three-channel grayscale image is created, and non-zero pixels of the binary mask are set to 255 in the first channel (arbitrarily chosen), effectively highlighting the cube.